11.3 Cost of preferred stock.  Kyle is raising funds for his company by selling preferred stock....

11.3

Cost of preferred stock.  Kyle is raising funds for his company by selling preferred stock. The preferred stock has a par value of $94and a dividend rate of 9.6​%. The stock is selling for $62.68 in the market. Kyle hires Wilson Investment Bankers to sell the preferred stock. Wilson charges a fee of   2​% on the sale of preferred stock. What is the cost of preferred stock for Kyle using the investment​banker?

What is the cost of preferred stock for Kyle using the investment​ banker?

____% (Round to two decimal places)
Answer #1

Dear student thankyou for using homeworklib.

Please find the answer below.

Statement showing computation:

Preferred dividend= 94*0.096= 9.024

Current price= 62.68*(1-0.02)= $61.4264

Cost of preferred stock= preferred dividend/current price

Cost of preferred stock= 9.024/61.4264= 14.69%

Cost of preferred stock= 14.69%